OpenStudy (anonymous):

mc002-1.jpg O is the center of the circle. Assume that lines that appear to be tangent are tangent. What is the value of x?

OpenStudy (anonymous):



OpenStudy (jdoe0001):

can we also have a screenshot maybe?

OpenStudy (jdoe0001):

http://phgeometryfinalexamreview.wikispaces.com/file/view/VERY_IMPORTANT.png/148266885/741x242/VERY_IMPORTANT.png the picture there uses a secant and a tangent but is pretty much the same thing really all you'd need are the INTERCEPTED CENTRAL ANGLES the central angle given there is 157 degrees what do you think is the other arc's central angle is?

OpenStudy (anonymous):

78.5?

OpenStudy (jdoe0001):

well, look at the circle keep in mind the circumference of the circle is 360 degrees

OpenStudy (anonymous):

101.5?

OpenStudy (jdoe0001):

well, the lines are touching the circle at the central angle of 157 degrees so the other bigger central angle will be 360 - 157 the rest of 360 once 157 is subtracted from it

OpenStudy (anonymous):

203

OpenStudy (jdoe0001):

yes, so the angle of "x", of the outside intersecting tangents, will be \(\bf \cfrac{\textit{far arc - near arc}}{2}\)

OpenStudy (anonymous):

203/2

OpenStudy (jdoe0001):

you're forgetting the "near arc", which is given

OpenStudy (anonymous):

so 203-157=46/2=23

OpenStudy (jdoe0001):

yes, \(\bf x = 23^o\)

OpenStudy (anonymous):

can you help me with another problem?

OpenStudy (jdoe0001):

I can try :)

OpenStudy (anonymous):

What is the value of x to the nearest tenth?

OpenStudy (jdoe0001):

 notice the line going across the circle is the diameter so if one side is 16, the other side is also 16, thus the radius is 16 now look at the triangle, using pythagorean theorem, what do you think is the missing vertical leg/side? \(\bf c^2 = a^2 + b^2 \implies c^2 - a^2 = b^2 \implies \sqrt{c^2-a^2} = b\)

OpenStudy (anonymous):

I don't understand how to do that.

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=AA6RfgP-AHU

OpenStudy (anonymous):

then its 9.6